Sand Castle Manufacturing produces concrete yard art. 2,600 concrete seahorses were produced in a recent production run. The run required 1,300 machine hours, and also required 6 "set-ups" of mixing equipment. Final inspection required 60 hours of inspection activity. Estimated overhead is estimated at $40 per machine hour, plus $2,800 per "set-up" and $30 per inspection hour. Direct materials and direct labor total $80 per seahorse.

(a) Apply activity-based costing and determine the amount assigned to a concrete seahorse.

A.

Amount assigned to a concrete seahorse

114.15

B.

Amount assigned to a concrete seahorse

125.15

C.

Amount assigned to a concrete seahorse

96.15

D.

Amount assigned to a concrete seahorse

107.15

E.

Amount assigned to a concrete seahorse

98.15
Answer #1
Machine hours activity 52000 =1300*40
Set-up activity 16800 =2800*6
Inspection activity 1800 =60*30
Overhead assigned 70600
Divide by Total concrete seahorses 2600
Overhead per concrete seahorse 27.15
Direct materials and direct labor 80
Overhead per concrete seahorse 27.15
Amount assigned to a concrete seahorse 107.15
Option D is correct